Given the balanced equation below for the highly exothermic reaction of aluminum



with oxygen to form aluminum oxide:



4Al(s) + 3O2(g)  2Al2O3(s) ∆H = - 1,670 kJ



a. What is the value of ∆H when 2 moles of aluminum react?



b. What is the value of ∆H when 10.8 g of aluminum reacts?



c. A certain reaction releases 6,881 kJ of heat. What mass of aluminum oxide was



produced in the reaction?

1
Expert's answer
2022-04-20T14:51:04-0400

4Al(s) + 3O2(g) ® 2Al2O3(s) ∆H = - 1,670 kJ

a. What is the value of ∆H when 2 moles of aluminum react?

2 moles Al correspond to the half of the given ΔH value, as 4 moles of Al is in the balanced equation:

-1,670 kJ / 2 = -835 kJ.

b. What is the value of ∆H when 10.8 g of aluminum reacts?

10.8 / 27 = 0.4 mol Al – is 10 times less of the value described in the balanced equation:

1,670 / 10 = 167 kJ.

c. A certain reaction releases 6,881 kJ of heat. What mass of aluminum oxide was produced in the reaction?

-6,881 / -1,670 = 4.12 times of the given value

4.12 * 4 mol = 16.48 mol

16.48 * 27 = 445 g Al